OpenStudy (anonymous):

evaluate [(21+6)-3²]/9*2

OpenStudy (anonymous):

Is it : \[\frac{(21+6)-3^2}{9\times2}\] ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Is it : \[\frac{(21+6)-3^2}9\times 2 ~~???\]